# Daejeon 34141, South Korea

### v1

@@ 4.1. $k$-convexity
We continue to assume that $\Gamma_{M}$ is hyperbolic. Let $\rho:\Gamma_{M}\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be a representation. We say that the representation $\rho$ is $k$-convex if there exist continuous maps $\zeta:\partial_{\infty}\Gamma_{M}\rightarrow\mathcal{G}_{k}({\mathbb{R}}^{n})$ and $\theta:\partial_{\infty}\Gamma_{M}\rightarrow\mathcal{G}_{n-k}({\mathbb{R}}^{n})$ such that the following hold:
@@
It follows that $\hat{\psi}_{t}$ induces in a flow on $E:={\mathbb{R}}^{n}_{\rho}=\hat{E}/\Gamma_{M}$.
@@
$\hat{Z}(x)=\zeta(x_{+})\hbox{ and }\hat{\Theta}(x)=\theta(x_{-})\hbox{ for }x\in\partial^{(2)}_{\infty}\hat{M}\times{\mathbb{R}}$
@@
Suppose that we further require for a $k$-dominated representation $\rho$ for $k\leq n/2$:
@@
Now, $\rho$ is also nkn-k-dominated. By Lemma 4.1, we obtain a new splitting of the bundle $E_{\rho}=Z^{\prime}\oplus\Theta^{\prime}$ over $\partial_{\infty}^{(2)}\hat{M}\times{\mathbb{R}}$ where $\dim Z^{\prime}=n-k$ and $\dim\Theta^{\prime}$ is $k$. Furthermore, $Z^{\prime}$ dominates $\Theta^{\prime}$. Then $Z$ is a subbundle of $Z^{\prime}$, and $\Theta^{\prime}$ is a subbundle of $\Theta$ by Proposition 2.1 of [5]. We now form bundles $Z,Z^{\prime}\cap\Theta,\Theta^{\prime}$ over $\partial_{\infty}^{(2)}\hat{M}\times{\mathbb{R}}$. The dominance property (1) of Definition 1.1 follows from those of $Z,\Theta$ and $Z^{\prime},\Theta^{\prime}$.
There are exponential expansions of $a_{i}(g)$ for $1\leq i\leq k$ by the partially hyperbolic condition of the premise. Now (iii)(a) of Definition 1.1 follows: Define $U_{p}(A)$ as the sum of eigenspaces of $\sqrt{AA^{*}}$ corresponding to $p$ largest eigenvalues. Define $S_{n-p}(A):=U_{n-p}(A^{-1})$. Since $\hat{M}/\Gamma_{M}$ is compact, Theorem 2.2 of [5] shows
@@ 4.2. Promoting $P$-Anosov representations to partially hyperbolic ones
We worked out with the following proposition with the help of Danciger and Stecker:
@@
      there exists a constant $C>0$ such that $|\log a_{i_{g}}(g)|<C$ for all $g\in\pi_{1}(M)$ where the index $1\geq i_{g}\geq n$ may depend on $g$.
@@
Let $B:=\mathcal{BC}_{Z}(\rho(G))$ denote the Benoist cone of $\rho(G)$ in $Z$. Let $S$ be a generating subset of $\rho(G)$. Then $B$ is a span of the joint spectrum $J(S)$ and 0 as stated immediately after Theorem 1.3 of [7]. The image $\iota_{Z}(B)$ is not convex in general.
Let $S\subset\rho(\pi_{1}(N))$ be a generating set. Under the condition (i), each vector $\vec{\lambda}(\rho(g)),g\in\pi_{1}(N)$ has a coordinate with $1$ as a value. By Theorem 1.6 of [7], we have a geometric convergence
$\vec{\lambda}_{Z}(S^{n})\rightarrow J(S)\hbox{ as }n\rightarrow\infty.$
Thus for each $x\in J(S)$, there is a sequence of elements of $x_{n}\in S^{n}$ converging to $x$. Since $B$ is the span of $J(S)$ and 0 in $a^{+}_{Z}$, we find that there must be an index $i$ where $\iota_{Z}(x)_{i}=0$ for every $x\in B$.
Under the condition (ii), for each element $x\in B$, there is a sequence $\{b_{i}\in S\}$ such that
$\lim_{n\rightarrow\infty}\log\vec{a}_{Z}(b_{1}\dots b_{n})=x.$
We find that there must be an index $i$ where $\iota_{Z}(b)=0$ for every $x\in B$.
@@
$\mathcal{I}(b)=\{i|\log\lambda_{i}(\iota_{Z}(b))=0\}\hbox{ for }b\in B.$
which is a consecutive set under the first premise. Let S([1,,n])S([1,...,n]) denote the collection of subsets of consecutive elements. We define an integral-interval function
$\mathcal{I}:B-\{0\}\rightarrow S([1,\dots,n])$
@@
(5) $\lim_{i\rightarrow\infty}\mathcal{I}(x_{i})\subset I(x)\hbox{ whenever }x_{i}\rightarrow x,x_{i},x\in B.$
Let $S$ denote a finite set of generators of $\rho(G)$ and their inverses. By Theorems 1.3 and 1.4 of [7], the sequence $\frac{1}{m}\log\vec{\lambda}_{Z}(S^{m})\subset J(S)$ geometrically converges to a compact convex set $J(S)\subset{\rm Cl}(\mathfrak{a}_{Z}^{+})$ as $m\rightarrow\infty$ (see Section 3.1 of [7]). Moreover, the cone spanned by 0 and $J(S)$ is the Benoist cone $B$. This implies that the set of directions of $\log\vec{\lambda}_{Z}(\rho(G))$ is dense in $B$. (See also [4]). Let $S_{n}$ denote the unit sphere in the Lie algebra $\mathfrak{a}_{n}$ of ${\mathsf{GL}}(n,{\mathbb{R}})$ with respect to the standard coordinates. We obtain
@@
for every $b\in B-\{O\}$ going to an element of $\iota_{Z}(B)\cap S_{n}$ for a constant $C>0$.
(II) Our first major step is to prove $\mathcal{I}(B-\{0\})\subset[k+1,n-k]$: Let $\iota_{n}$ denote the opposite involution of $\mathfrak{a}_{n}$. First, suppose that the rank of $Z$ is $\geq 2$, and hence $\dim\mathfrak{a}_{Z}^{+}\geq 2$.
@@
$\Delta_{i,j}\subset\mathfrak{a}_{n}\hbox{ for }i\neq j,i,j=1,\dots,n,$
is a subspace of the maximal abelian algebra $\mathfrak{a}_{n}$ given by $\log\lambda_{i}-\log\lambda_{j}=0$ for some indices $i$ and $j$. The inverse images under $\iota_{Z}$ in ${\rm Cl}(\mathfrak{a}_{Z}^{+})$ of the diagonal subspaces of $\mathfrak{a}$ may meet $B$. We let $\mathcal{I}_{B}$ the maximal collection of indices $(i,j),i<j$, where $\iota_{Z}^{-1}(\Delta_{i,j})$ contains $B$, which may be empty. Let $\mathcal{I}^{\prime}_{B}$ the collection of indices $(k,l)$, $k<l$, of $\Delta_{k,l}$ not in $\mathcal{I}_{B}$. We call the closures of the components of
@@
the generic flat domains of $B$. There are finitely many of these to be denoted $B_{1},\dots,B_{m}$ which are convex domains in $B$.
@@
Heuristically speaking: the worms leave “traces”.
We now aim to prove that the following never occurs –(*):
There exists a pair of elements $b_{1},b_{2}\in B$ with $b_{1}\in B_{p}^{o}$ and $b_{2}\in B_{q}^{o}$ for some $p,q$ where $\mathcal{I}(b_{1})$ has an element $\leq k$ and $\mathcal{I}(b_{2})$ has an element >k>k.
@@
(7) $\mathcal{I}(B_{l_{p}}^{o})\cup\mathcal{I}(B_{l_{p+1}}^{o})\subset\mathcal{I}(x)\hbox{ for }x\in{\rm Cl}(B_{l_{p}})\cap{\rm Cl}(B_{l_{p+1}})-\{O\}.$
Considering $\mathcal{I}(B_{l_{j}}^{o})$, $j=1,\dots,m$, and the $\mathcal{I}$-values of the intersections of adjacent generic flat domains, we have $\mathcal{I}(b^{\prime})\ni k,k+1$ for some nonzero element $b^{\prime}$ of $B$ by the connectedness of the $\mathcal{I}$-values as we change generic flat domains to adjacent ones according to (7). Hence, we obtain
@@
  • for all $b\in B$ in the interiors of generic flat domains, every element of $\mathcal{I}(b)$ is $\leq k$ or
  • for all $b\in B$ in the interiors of generic flat domains, every element of $\mathcal{I}(b)$ is >k>k.
Note that the opposition involution $\iota_{n}$ sends $\iota_{Z}(B)$ to itself since the inverse map $g\rightarrow g^{-1}$ preserves $Z$ and the set of eigenvalues of $g$ is sent to the set of the eigenvalues of $g^{-1}$. In the first case, by the opposition involution $\iota_{n}$ preserving $G$, $\mathcal{I}(\iota_{n}(b))$ for $b\in B$ contains an element $\geq n-k+1$. This again contradicts the above. Hence, we conclude $\mathcal{I}(b)\subset[k+1,\dots,n]$ for all $b\in B$ in the interiors of generic flat domains. Acting by the opposition involution and using by similar arguments, we obtain $\mathcal{I}(b)\subset[k+1,\dots,n-k]$ for all $b\in B$ in the interiors of generic flat domains.
Furthermore, we can show that $\mathcal{I}(b)\subset[k+1,n-k]$ for all $b\in B-\{0\}$ since otherwise, we have $\mathcal{I}(b^{\prime})\ni k,k+1$ for some $b^{\prime}\neq 0$ by a similar reason to the above applied to lower-dimensional strata. This is a contradiction as before. In addition, the argument shows $k+1,n-k]\neq\emptyset$ and $k<n/2$.
@@
$\mathcal{I}(B-\{0\})=[i,\dots,n-i+1]\hbox{ for some index }i,0<i\leq n/2.$
Now by (6), we must have $k<i$. This implies $\mathcal{I}(B-\{0\})\subset[k+1,n-k]$.
@@
$\left\{\frac{\log\vec{a}(\rho(g_{l})))}{w(g_{l})}\right\}$
have limit points only in $\iota_{Z}(B)$ by Theorem 1.3 of [7]. By the conclusion of Step (II), for any sequence $\{g_{l}\}$ in $G$,
$\left\{\frac{\min\{|\log a_{i}(\rho(g_{l}))|\}_{i=k+1,\dots n-k}}{w(g_{l})}\right\}\rightarrow 0,$
@@
Suppose that $\rho$ is $P$-Anosov in the sense of [24]. Hence, $\Gamma:=\rho(\pi_{1}(N))$ is word-hyperbolic by [24].
First, suppose that the Zariski closure of the image of the linear holonomy group is reductive. By Lemma 3.1, $\rho$ is a $k$-dominated representation of the index $k$. We take a finite-index normal subgroup $\Gamma^{\prime}$ of $\Gamma$. Assume that the Zariski closure $Z$ of $\rho(\Gamma^{\prime})$ is a connected reductive Lie group. By Proposition 4.3, $\rho|\Gamma^{\prime}$ is partially hyperbolic in the singular-value sense with the index $k$, $k<n/2$. So is $\rho$ since $\Gamma$ is a finite-index extension of $\Gamma^{\prime}$. By Theorem 4.2, $\rho$ is a partially hyperbolic representation in the bundle sense with the index $k$.
Now, we drop the reductive condition of the Zariski closure. Mainly, we have to prove the expanding and contracting properties in the subbundles. The semisimplification $\rho^{ss}$ is $P$-Anosov by Proposition 1.8 of [24]. Since conjugations do not change eigenvalues, $\rho^{ss}(g)$ for each $g\in\Gamma$ has an eigenvalue equal to $1$ by Theorem 5.1. (See also the proof of Lemma 2.40 of [24].) By Proposition 4.3 and Theorem 4.2, $\rho^{ss}$ is a partially hyperbolic representation in the bundle sense with the index $k$, $k<n/2$. By Proposition 1.3 of [24], $\rho$ is $P$-Anosov with respect to a continuous Anosov map $\zeta:\partial_{\infty}\Gamma\rightarrow{\mathcal{F}}$ for the flag variety $\mathcal{F}:={\mathsf{GL}}(n,{\mathbb{R}})/P$.
We can conjugate $\rho$ arbitrarily close to $\rho^{ss}$ in ${\rm Hom}(\Gamma,{\mathsf{GL}}(n,{\mathbb{R}}))$ since the orbit of $\rho^{ss}$ is the closed orbit in the closure of the orbit of $\rho$ under the conjugation action. (See Section 2.5.4 of [24]. ) Let $\zeta_{i}:\partial_{\infty}\Gamma\rightarrow{\mathcal{F}}$ denote the Anosov map for $\rho_{i}$ conjugate to $\rho$ converging to $\rho^{ss}$. Theorem 5.13 of [25] generalizes to ${\mathsf{GL}}(n,{\mathbb{R}})$ since we can multiply a representation by a homomorphism $\Gamma\rightarrow{\mathbb{R}}^{+}$ not changing the $P$-Anosov property to ${\mathsf{SL}}_{\pm}(n,{\mathbb{R}})$-representations. (See Section 2.5.3 of [24]). Hence, Theorem 5.13 of [25] implies that $\zeta_{i}$ converges to $\zeta^{ss}:\partial_{\infty}\Gamma\rightarrow{\mathcal{F}}$ the map for $\rho^{ss}$ in the $C^{0}$-topology.
@@
  • Let $\hat{\mathbb{V}}^{ss+},\hat{\mathbb{V}}^{ss0}$, and $\hat{\mathbb{V}}^{ss-}$ denote the cover of ${\mathbb{V}}^{ss+},{\mathbb{V}}^{ss0}$, and ${\mathbb{V}}^{ss-}$.
  • Let $\hat{\mathbb{V}}^{+}_{i},\hat{\mathbb{V}}^{0}_{i}$, and $\hat{\mathbb{V}}^{-}_{i}$ denote the cover of ${\mathbb{V}}^{+}_{i},{\mathbb{V}}^{0}_{i}$, and ${\mathbb{V}}^{-}_{i}$.
@@
(9) $\hat{\mathbb{V}}^{+}_{i}(x)\rightarrow\hat{\mathbb{V}}^{ss+}(x),\hat{\mathbb{V}}^{0}_{i}(x)\rightarrow\hat{\mathbb{V}}^{ss0}(x),\hbox{ and }\hat{\mathbb{V}}^{-}_{i}(x)\rightarrow\hat{\mathbb{V}}^{ss-}(x)\hbox{ as }i\rightarrow\infty$
@@
The convergence is also uniform over every compact subset of $\mathbf{UC}\hat{M}$ since these vector spaces depends only on the endpoints of quasi-geodesic though $x$ and $\hat{M}$ is quasi-isometric to $\partial^{2}_{\infty}\Gamma_{M}\times{\mathbb{R}}$.
@@
Hence for sufficiently large $i$, the flow $\Phi_{i}$ on ${\mathbb{R}}^{n}_{\rho_{i}}$ satisfies the corresponding properties as well as the domination properties for $\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|_{{\mathbb{R}}^{n}_{\rho_{i}}}$ by (9). Hence, $\Phi_{i}$ is partially hyperbolic with the index $k$.
For converse, suppose that $\rho$ is partially hyperbolic with the index $k$. Proposition 4.5 of [5] shows that $\rho$ is $k$-dominated. Lemma 3.1 shows that $\rho$ is $P_{\theta}$-Anosov for $k\in\theta$. Also, $k<n/2$ since the neutral bundle exists. ∎

### v2

@@ 4.1. $k$-convexity
We continue to assume that $\Gamma_{M}$ is word-hyperbolic. Let $\rho:\Gamma_{M}\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be a representation. We say that the representation $\rho$ is $k$-convex if there exist continuous maps $\zeta:\partial_{\infty}\Gamma_{M}\rightarrow\mathcal{G}_{k}({\mathbb{R}}^{n})$ and $\theta:\partial_{\infty}\Gamma_{M}\rightarrow\mathcal{G}_{n-k}({\mathbb{R}}^{n})$ such that the following hold:
@@
It follows that $\hat{\psi}_{t}$ induces a flow on $E_{\rho}:={\mathbb{R}}^{n}_{\rho}=\hat{E}/\Gamma_{M}$.
@@
$\hat{Z}(x)=\zeta(x_{+})\text{ and }\hat{\Theta}(x)=\theta(x_{-})\text{ for }x\in\partial^{(2)}_{\infty}\hat{M}\times{\mathbb{R}}$
@@
Suppose that we further require for a $k$-dominated representation $\rho$ where $k\leq n/2$:
@@
Now, $\rho$ is also (nk)(n-k)-dominated. By Lemma 4.1, we obtain a new splitting of the bundle $E_{\rho}=Z^{\prime}\oplus\Theta^{\prime}$ over $\partial_{\infty}^{(2)}\hat{M}\times{\mathbb{R}}$ where $\dim Z^{\prime}=n-k$ and $\dim\Theta^{\prime}$ is $k$. Furthermore, $Z^{\prime}$ dominates $\Theta^{\prime}$. Then $Z$ is a subbundle of $Z^{\prime}$, and $\Theta^{\prime}$ is a subbundle of $\Theta$ by Proposition 2.1 of [5]. We now form bundles $Z,Z^{\prime}\cap\Theta,\Theta^{\prime}$ over $\partial_{\infty}^{(2)}\hat{M}\times{\mathbb{R}}$. The dominance property (1) of Definition 1.1 follows from those of $Z,\Theta$ and $Z^{\prime},\Theta^{\prime}$.
There are exponential expansions of $a_{i}(g)$ for $1\leq i\leq k$ by the partially hyperbolic condition of the premise. Now (iii)(a) of Definition 1.1 follows: Define $U_{p}(A)$ as the sum of eigenspaces of $\sqrt{AA^{*}}$ corresponding to the $p$ largest eigenvalues. Define $S_{n-p}(A):=U_{n-p}(A^{-1})$. Since $\hat{M}/\Gamma_{M}$ is compact, Theorem 2.2 of [5] shows
@@ 4.2. Promoting $P$-Anosov representations to partially hyperbolic ones
By a cone of a subset $J$ of a vector space $V$, we mean the subset of nonnegative multiples of points of $J$. We denote it by $C(J)$. This is not necessarily convex..
@@
We worked out the following proposition with the help of Danciger and Stecker:
@@
      For any sequence $g_{i}\in\pi_{1}(N)$, if $w(g_{i})\rightarrow\infty$, then $\frac{|\log a_{j_{g_{i}}}(\rho(g_{i}))|}{w(g_{i})}\rightarrow 0$ for some choices of the indices $j_{g_{i}}\in\{1,\dots,n\}$ depending on $g_{i}$.
@@
Let $B:=\mathcal{BC}_{Z}(\rho(G))$ denote the Benoist cone of $\rho(G)$ in $Z$. Let $S$ be a generating subset of $\rho(G)$. Then $B$ is a cone of the joint spectrum $J(S)$ as stated immediately after Theorem 1.3 of [7]. That is, $B=C(J(S))$. The image $\iota_{Z}(B)$ is not convex in general.
Let $S\subset\rho(G)$ be a generating set. The condition (ii) implies (i) by Lemma 2.11 of [7]. Under the condition (i), each vector $\log\vec{\lambda}(\rho(g)),g\in G$ has a coordinate with the value 0. By Theorem 1.2 of [7], we have
${\rm Cl}\left(\bigcup_{m=1}^{\infty}\frac{1}{m}\log\vec{\lambda}_{Z}(S^{m})\right)=J(S)$
Thus for each $x\in J(S)$, there is a sequence of elements of $x_{j}\in\frac{1}{m_{j}}\log\lambda_{Z}(S^{m_{j}})$ converging to $x$. Since $B$ is the cone of $J(S)$ in $a^{+}_{Z}$, we find that there must be an index $i$ where $\iota_{Z}(x)_{i}=0$ for every $x\in B$.
@@
$\mathcal{I}(b)=\{i\mid\log\lambda_{i}(\iota_{Z}(b))=0\}\text{ for }b\in B.$
which is a consecutive set under the first premise. Let S([1,,n])S([1,...,n]) denote the collection of subsets of consecutive elements.
We define an integral-interval function,
$\mathcal{I}:B\setminus\{O\}\rightarrow S([1,\dots,n])$
@@
(5) $\lim_{i\rightarrow\infty}\mathcal{I}(x_{i})\subset\mathcal{I}(x)\text{ whenever }x_{i}\rightarrow x,x_{i},x\in B.$
Let $S$ denote a finite set of generators of $\rho(G)$ and their inverses. By Theorem 1.3 of [7], the sequence $\frac{1}{m}\log\vec{\lambda}_{Z}(S^{m})\subset J(S)$ geometrically converges to the compact convex set $J(S)\subset{\rm Cl}(\mathfrak{a}_{Z}^{+})$ as $m\rightarrow\infty$ (see Section 3.1 of [7]). Moreover, the cone of $J(S)$ is the Benoist cone $B$. This implies that the set of directions of $\log\vec{\lambda}_{Z}(\rho(G))$ is dense in that of $B$. (See also [4]). Let $S_{n}$ denote the unit sphere in the Lie algebra $\mathfrak{a}_{n}$ of ${\mathsf{GL}}(n,{\mathbb{R}})$ with respect to the standard coordinates. We obtain
@@
for every $b\in B\setminus\{O\}$ going to an element of $\iota_{Z}(B)\cap S_{n}$ for a constant $C>0$.
(II) Our first major step is to prove $\mathcal{I}(B\setminus\{O\})\subset[k+1,n-k]$: Let $\iota_{n}$ denote the opposite involution of $\mathfrak{a}_{n}$. First, suppose that the rank of $Z$ is $\geq 2$, and hence $\dim\mathfrak{a}_{Z}^{+}\geq 2$.
Recall from the above of this proposition that $B\setminus\{O\}$ is path-connected.
@@
$\Delta_{i,j}\subset\mathfrak{a}_{n}\text{ for }i\neq j,i,j=1,\dots,n,$
is a subspace of the maximal abelian algebra $\mathfrak{a}_{n}$ given by $\log\lambda_{i}-\log\lambda_{j}=0$ for some indices $i$ and $j$. The inverse images under $\iota_{Z}$ in ${\rm Cl}(\mathfrak{a}_{Z}^{+})$ of the diagonal subspaces of $\mathfrak{a}_{n}$ may meet $B$.
Hueristically speaking, under $\iota_{Z}$, $B$ is mapped piecewise linearly in $\mathfrak{a}_{n}$ and when it meets $\Delta_{i,j}$ for some $i,j$, it continues in the subspace obtained by applying a reflection about $\Delta_{i,j}$.
Let $\mathcal{I}^{\prime}_{B}$ the collection of indices $(k,l)$, $k<l$, of $\Delta_{k,l}$ so that $B$ meets $\Delta_{k,l}$ transversally. We call the closures of the components of
@@
the generic flat cones of $B$. There are finitely many of these to be denoted $B_{1},\dots,B_{m}$ which are convex cones in $B$. These are mapped linearly inside the positive Weyl chamber of $\mathfrak{a}_{n}$.
@@
Heuristically speaking, worms leave “traces”.
We now aim to prove that the following never occurs:
There exists a pair of elements $b_{1},b_{2}\in B$ with $b_{1}\in B_{p}^{o}$ and $b_{2}\in B_{q}^{o}$ for some $p,q$ where $\mathcal{I}(b_{1})$ has an element $\leq k$ and $\mathcal{I}(b_{2})$ has an element >k>k—(*).
@@
(7) $\mathcal{I}(B_{l_{p}}^{o})\cup\mathcal{I}(B_{l_{p+1}}^{o})\subset\mathcal{I}(x)\text{ for }x\in{\rm Cl}(B_{l_{p}})\cap{\rm Cl}(B_{l_{p+1}})\setminus\{O\}.$
Considering $\mathcal{I}(B_{l_{j}}^{o})$, $j=1,\dots,m$, and the $\mathcal{I}$-values of the intersections of adjacent generic flat cones, we have $\mathcal{I}(b^{\prime})\ni k,k+1$ for some nonzero element $b^{\prime}$ of $B$ by the connectedness of the $\mathcal{I}$-values as we change generic flat cones to adjacent ones according to (7). Hence, we obtain
@@
  • for all $b\in B$ in the interiors of generic flat cones, every element of $\mathcal{I}(b)$ is $\leq k$ or
  • for all $b\in B$ in the interiors of generic flat cones, every element of $\mathcal{I}(b)$ is >k>k.
Note that the opposition involution $\iota_{n}$ sends $\iota_{Z}(B)$ to itself since the inverse map $g\rightarrow g^{-1}$ preserves $Z$ and the set of eigenvalues of $g$ is sent to the set of the eigenvalues of $g^{-1}$. In the first case, by the opposition involution $\iota_{n}$ preserving $G$, $\mathcal{I}(\iota_{n}(b))$ for $b\in B$ contains an element $\geq n-k+1$. This again contradicts the above. Hence, we conclude $\mathcal{I}(b)\subset[k+1,\dots,n]$ for all $b\in B$ in the interiors of generic flat cones. Acting by the opposition involution and using similar arguments, we obtain $\mathcal{I}(b)\subset[k+1,\dots,n-k]$ for all $b\in B$ in the interiors of generic flat cones.
Furthermore, we can show that $\mathcal{I}(b)\subset[k+1,n-k]$ for all $b\in B\setminus\{O\}$ since otherwise, we have $\mathcal{I}(b^{\prime})\ni k,k+1$ for some $b^{\prime}\neq O$ by a similar reason to the above applied to lower-dimensional strata. This is a contradiction as before. In addition, the argument shows $k+1,n-k]\neq\emptyset$ and $k<n/2$.
@@
$\mathcal{I}(B\setminus\{O\})=[i,\dots,n-i+1]\hbox{ for some index }i,0<i\leq n/2.$
Now by (6), we must have $k<i$. This implies $\mathcal{I}(B\setminus\{O\})\subset[k+1,n-k]$.
@@
$\left\{\frac{\log\vec{a}(\rho(g_{l}))}{w(g_{l})}\right\}$
has limit points only in $\iota_{Z}(B)$ by Theorem 1.3 of [7]. By the conclusion of Step (II), for any sequence $\{g_{l}\}$ in $G$,
$\frac{\min_{i=k+1,\dots,n-k}|\log a_{i}(\rho(g_{l}))|}{w(g_{l})}\rightarrow 0,$
@@
Suppose that $\rho$ is $P_{\theta}$-Anosov in the sense of [24] where $k\in\theta$ and $\theta^{\ast}=\theta$. Hence, $\Gamma:=\rho(\pi_{1}(N))$ is word-hyperbolic by the definition in [24].
First, suppose that the Zariski closure of the image of the linear holonomy group is reductive. By Lemma 3.1, $\rho$ is a $k$-dominated representation of the index $k$. We take a finite-index normal subgroup $\Gamma^{\prime}$ of $\Gamma$ so that the Zariski closure $Z$ of $\rho(\Gamma^{\prime})$ is a connected reductive Lie group. By Proposition 4.3, $\rho|\Gamma^{\prime}$ is partially hyperbolic in the singular-value sense with the index $k$, $k<n/2$. So is $\rho$ since $\Gamma$ is a finite-index extension of $\Gamma^{\prime}$. By Theorem 4.2, $\rho$ is a partially hyperbolic representation in the bundle sense with the index $k$.
Now, we drop the reductive condition of the Zariski closure. Mainly, we have to prove the expanding and contracting properties in the subbundles. From now on, $P:=P_{\theta}$. The semisimplification $\rho^{ss}$ is $P$-Anosov by Proposition 1.8 of [24], and the Zariski closure of the image of $\rho$ is reductive. We can conjugate $\rho$ arbitrarily close to $\rho^{ss}$ in ${\rm Hom}(\Gamma,{\mathsf{GL}}(n,{\mathbb{R}}))$ since the orbit of $\rho^{ss}$ is the closed orbit in the closure of the orbit of $\rho$ under the conjugation action. (See Section 2.5.4 of [24]. ) We can assume that $\rho$ has the property (i) by Lemma 2.11 of [7]. Since eigenvalues do not change under conjugation, $\rho^{ss}(g)$ for each $g\in\Gamma$ has an eigenvalue equal to $1$. Hence (i) holds for $\rho^{ss}$. (See also the proof of Lemma 2.40 of [24].) By Proposition 4.3 and Theorem 4.2, $\rho^{ss}$ is a partially hyperbolic representation in the bundle sense with the index $k$, $k<n/2$. By Proposition 1.3 of [24], $\rho$ is $P$-Anosov with respect to a continuous Anosov map $\zeta:\partial_{\infty}\Gamma\rightarrow{\mathcal{F}}$ for the flag variety $\mathcal{F}:={\mathsf{GL}}(n,{\mathbb{R}})/P$.
Let $\zeta_{i}:\partial_{\infty}\Gamma\rightarrow{\mathcal{F}}$ denote the Anosov map for $\rho_{i}$ conjugate to $\rho$ converging to $\rho^{ss}$. Theorem 5.13 of [25] generalizes to ${\mathsf{GL}}(n,{\mathbb{R}})$ since we can multiply a representation by a homomorphism $\Gamma\rightarrow{\mathbb{R}}^{+}$ not changing the $P$-Anosov property to ${\mathsf{SL}}_{\pm}(n,{\mathbb{R}})$-representations. (See Section 2.5.3 of [24]). Hence, Theorem 5.13 of [25] implies that $\zeta_{i}$ converges to $\zeta^{ss}:\partial_{\infty}\Gamma\rightarrow{\mathcal{F}}$ the map for $\rho^{ss}$ in the $C^{0}$-topology.
@@
  • Let $\hat{\mathbb{V}}^{ss+},\hat{\mathbb{V}}^{ss0}$, and $\hat{\mathbb{V}}^{ss-}$ denote the cover of ${\mathbb{V}}^{ss+},{\mathbb{V}}^{ss0}$, and ${\mathbb{V}}^{ss-}$ respectively.
  • Let $\hat{\mathbb{V}}^{+}_{i},\hat{\mathbb{V}}^{0}_{i}$, and $\hat{\mathbb{V}}^{-}_{i}$ denote the cover of ${\mathbb{V}}^{+}_{i},{\mathbb{V}}^{0}_{i}$, and ${\mathbb{V}}^{-}_{i}$ respectively.
@@
(9) $\hat{\mathbb{V}}^{+}_{i}(x)\rightarrow\hat{\mathbb{V}}^{ss+}(x),\hat{\mathbb{V}}^{0}_{i}(x)\rightarrow\hat{\mathbb{V}}^{ss0}(x),\text{ and }\hat{\mathbb{V}}^{-}_{i}(x)\rightarrow\hat{\mathbb{V}}^{ss-}(x)\text{ as }i\rightarrow\infty$
@@
The convergence is also uniform over every compact subset of $\mathbf{UC}\hat{M}$ since these vector spaces depend only on the endpoints of quasi-geodesics through $x$ and $\hat{M}$ is quasi-isometric to $\partial^{(2)}_{\infty}\Gamma_{M}\times{\mathbb{R}}$.
@@
Hence, for sufficiently large $i$, the flow $\Phi_{i}$ on ${\mathbb{R}}^{n}_{\rho_{i}}$ satisfies the corresponding properties as well as the domination properties for $\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|_{{\mathbb{R}}^{n}_{\rho_{i}}}$ by (9). Hence, $\Phi_{i}$ is partially hyperbolic with the index $k$.
Conversely, suppose that $\rho$ is partially hyperbolic with the index $k$. Proposition 4.6 of [5] shows that $\rho$ is $k$-dominated. Lemma 3.1 (or Proposition 4.9 of [5]) shows that $\rho$ is $P_{\theta}$-Anosov for $k\in\theta$. Also, $k<n/2$ since the neutral bundle exists. ∎